# Characterizing the cage state of glassy systems and its sensitivity to frozen boundaries

### v1

@@ III.2 Structure of the cage state
Next, we examine the extent to which the measured CSs differ structurally across various cavity sizes. From this point on, we therefore no longer use the correlation with the dynamic propensity to evaluate the CS, but instead focus solely on its structural properties. When comparing the CS measured in a cavity to the CS of the unfrozen system, we note that the observed differences can be attributed to two main factors: boundary effects associated with the cavity, and thermodynamic noise arising from undersampling of the cage, which is present in both the frozen and unfrozen systems. Since we are only interested in the influence of the boundaries, we want to isolate this effect. In order to do so, we measure the unfrozen CS twice, denoted by $\{\mathbf{r}^{\text{cage,1}}\}$, $\{\mathbf{r}^{\text{cage,2}}\}$, and compute the difference between those two CSs, $\{\mathbf{r}^{\text{cage,1}}-\mathbf{r}^{\text{cage,2}}\}$. This difference serves as a reference for the noise due to undersampling. We can then define a measure $\Gamma$, which captures the difference between the cavity CS and the unfrozen CS that can be attributed to boundary effects,
@@
where the sum runs over all particles under consideration. When $\Gamma\approx 0$, the difference between the cavity cage center and the unfrozen cage centers is mainly due to thermodynamic noise, while a significant deviation from zero indicates boundary-induced differences. Note that we checked that when $R_{f}\rightarrow\infty$, $\Gamma$ is approximately zero. In Fig. 5 we plot $\Gamma$ as a function of freezing radius for six packing fractions. The results show that even for the lower packing fractions, at $R_{f}=7\sigma_{A}$, boundary conditions influence the cage center within the cavity. However, it is clear that at higher packing fractions, the range over which the boundary conditions influence the cage state is significantly higher. Note that this enhanced sensitivity of the structure to boundary conditions as the system becomes more glassy has also been observed in point-to-set functions, see e.g. Ref. biroli2008thermodynamic, , where it was attributed to a growing amorphous structural length scale. Although our results do not allow for a quantitative conclusion, they too suggest the presence of a growing amorphous structural length scale associated with the structure observed during the caging regime.
Finally, we study the positions of particles within their cages in more detail. Specifically, during the cage state measurements, we track the non-averaged displacement vectors of individual particles, denoted as $\delta\mathbf{r}^{\text{cage}}$. These vectors connect particle’s instantaneous positions to the associated cage center. For a specific particle, the collection of these vectors thus captures the fluctuations of the particle around its cage center. To quantify these fluctuations we compute the average standard deviation of these vectors
$SD=\frac{1}{N_{m}}\sum_{i}\sqrt{\langle\delta\mathbf{r}_{i}^{\text{cage}}\cdot\delta\mathbf{r}_{i}^{\text{cage}}\rangle-\langle\delta\mathbf{r}_{i}^{\text{cage}}\rangle^{2}},$
where the sum runs over $N_{m}$ particles. Note that by definition $\langle\delta\mathbf{r}_{i}^{\text{cage}}\rangle=\mathbf{0}$, such that the second term under the square root vanishes. Fig. 6 shows the standard deviations as a function of freezing radius for the six packing fractions considered. As expected, the standard deviation is higher at lower packing fractions, reflecting the weaker caging of particles. Additionally we find that the standard deviation increases monotonically with increasing freezing radius, indicating that the particles can fluctuate more as the cavity size increases. Notably, the standard deviation plateaus earlier for lower packing fractions. This suggests that, as the packing fraction increases, the freezing of particles increasingly affects not only the average positions of particles within their cages, but also the extent of their fluctuations.
@@ IV Conclusion
Overall, we found that the correlation between the CS and the dynamic propensity increases significantly with increasing packing fraction, indicating that the CS becomes better defined as the system becomes more glassy. Furthermore, we found that boundary conditions at large distances strongly affect this correlation – an effect that becomes more pronounced as the system approaches the glassy regime. This observed collective nature of the CS may help explain why machine learning techniques that are based solely on initial positions struggle to accurately predict particle dynamics during the caging regime. Our results suggest that, at high packing fractions, the structural information relevant to dynamical predictions is influenced by particles at large distances, implying that ML strategies must account for these non-local, collective effects in order to make good dynamical predictions.
@@
These findings further establish the CS as a valuable structural quantity for capturing the collective nature of the caging regime and for probing the dynamics that govern particle dynamics during those timescales.
@@ V Acknowledgments
All relevant data are available upon reasonable request.

### v2

@@ III.2 Structure of the cage state
Next, we examine the extent to which the measured CSs differ structurally across various cavity sizes. From this point on, we therefore no longer use the correlation with the dynamic propensity to evaluate the CS, but instead focus solely on its structural properties. When comparing the CS measured in a cavity to the CS of the unfrozen system, we note that the observed differences can be attributed to two main factors: boundary effects associated with the cavity, and thermodynamic noise arising from undersampling of the cage which is present in both the frozen and unfrozen systems. Since we are only interested in the influence of the boundaries, we want to isolate this last effect. In order to do so, we perform two independent simulations of the unfrozen system, generating two separate unfrozen CSs based on the same initial configuration, denoted by $\{\mathbf{r}^{\text{cage,1}}\}$, $\{\mathbf{r}^{\text{cage,2}}\}$, and compute the difference between those two CSs, $\{\mathbf{r}^{\text{cage,1}}-\mathbf{r}^{\text{cage,2}}\}$. This difference serves as a reference for the noise due to undersampling. We can then define a measure $\Gamma$, which captures the difference between the cavity CS and the unfrozen CS that can be attributed to boundary effects,
@@
where the sum runs over all particles under consideration. When $\Gamma\approx 0$, the difference between the cavity cage center and the unfrozen cage centers is mainly due to thermodynamic noise, while a significant deviation from zero indicates boundary-induced differences. Note that we checked that when $R_{f}\rightarrow\infty$, $\Gamma$ is approximately zero. In Fig. 5 we plot $\Gamma$ as a function of freezing radius for six packing fractions. The results show that at $R_{f}=7\sigma_{A}$, boundary conditions influence the cage center within the cavity even at the lower packing fractions. However, it is clear that at higher packing fractions, the range over which the boundary conditions influence the cage state is significantly higher. Note that this enhanced sensitivity of the structure to boundary conditions as the system becomes more glassy has also been observed in point-to-set functions, see e.g. Ref. biroli2008thermodynamic, , where it was attributed to a growing amorphous structural length scale. Although our results do not allow for a quantitative conclusion, they too suggest the presence of a growing amorphous structural length scale associated with the structure observed during the caging regime.
Finally, we study the positions of particles within their cages in more detail. Specifically, during the cage state measurements, we track the non-averaged displacement vectors of individual particles, denoted as $\delta\mathbf{r}^{\text{cage}}$. These vectors connect particle’s instantaneous positions to the associated cage centers. For a specific particle, the collection of these vectors thus captures the fluctuations of the particle around its cage center. To quantify these fluctuations we compute the average standard deviation of these vectors
$SD=\frac{1}{N_{m}}\sum_{i}\sqrt{\langle\delta\mathbf{r}_{i}^{\text{cage}}\cdot\delta\mathbf{r}_{i}^{\text{cage}}\rangle-\langle\delta\mathbf{r}_{i}^{\text{cage}}\rangle^{2}},$ (5)
where the sum runs over $N_{m}$ particles. Note that by definition $\langle\delta\mathbf{r}_{i}^{\text{cage}}\rangle=\mathbf{0}$, such that the second term under the square root vanishes. Fig. 6 shows the standard deviations as a function of freezing radius for the six packing fractions considered. As expected, the standard deviation is higher at lower packing fractions, reflecting the weaker caging of particles. Additionally, we find that the standard deviation increases monotonically with increasing freezing radius, indicating that the particles can fluctuate more as the cavity size increases. Notably, the standard deviation plateaus earlier for lower packing fractions. This suggests that, as the packing fraction increases, the freezing of particles increasingly affects not only the average positions of particles within their cages, but also the extent of their fluctuations.
@@ IV Conclusion
Overall, we found that the correlation between the CS and the dynamic propensity increases significantly with increasing packing fraction, indicating that the CS becomes better defined as the system becomes more glassy. In addition, we observed that with decreasing packing fraction, the choice of the restriction function for confining particles to their cages has more influence on the CS. At lower packing fractions, restricting particles to their approximate Voronoi cell led to significantly lower correlations with the propensity than the spherical restriction method. This difference can likely be attributed to the fact that, as the system becomes more liquid, a particle’s Voronoi cell, which is strictly larger than the restricted spherical volume, mimics the cage less accurately. We expect this finding to be generalizable to other glass formers, meaning that the spherical restriction method is preferred also in those systems.
Furthermore, we found that the correlation between the propensity and the CS measured in an unfrozen cavity is strongly affected by the boundary conditions at large distances – an effect that becomes more pronounced as the system approaches the glassy regime. The observed collective nature of the CS may help explain why machine learning techniques that are based solely on initial positions struggle to accurately predict particle dynamics during the caging regime: Our results suggest that, at high packing fractions, the structural information relevant to dynamical predictions is influenced by particles at large distances, implying that ML strategies must account for these non-local, collective effects in order to make good dynamical predictions.
@@
These findings further establish the CS as a valuable structural quantity for capturing the collective nature of the caging regime and for probing the dynamics that govern particle dynamics during those timescales. Note that to date the CS has been used in ML studies. However, we expect it to be useful more generally for measuring quantities that depend on either the initial or inherent state, such as the local potential energy, local free volume, or Voronoi-based metrics richard2020predicting ; jung2023predicting .
@@ V Acknowledgments
All simulation codes needed to reproduce the data, as well as all relevant simulation data and notebooks to analyze the data and generate the figures is published as a data package in Ref. alkemade_2025_16947949, .